# Coarse-to-Fine Grounded Memory for LLM Agent Planning

### v1

@@ 1 Introduction
Figure 1: Conceptual overview of CFGM. During training, our method first extracts coarse-grained focus points to guide experience collection, then distills actionable tips from these experiences. At inference time, it retrieves relevant experiences and tips for planning. When encountering anomalies, the system identifies fine-grained key details for adaptive self-QA and plan adjustment.
@@
The remarkable progress of Large Language Models (LLMs) has ignited intense interest in their deployment as autonomous agents capable of tackling complex, long-horizon tasks within interactive environments (Wang et al., 2023; Xi et al., 2023; Gur et al., 2024; Wang et al., 2024a). These LLM-based agents leverage extensive world knowledge (Zhao et al., 2023) and reasoning capabilities to formulate policies (Yao et al., 2023a). However, significant challenges hinder their practical effectiveness. Agents often suffer from inefficient exploration, yielding sparse or irrelevant data (Arumugam and Griffiths, 2025; Zhang et al., 2025b). They also struggle with reasoning derailment in partially observable environments, where incomplete observations cause cascading errors (Kim et al., 2024). Existing methods attempt to mitigate these issues using memory mechanisms, such as storing offline experiences and insights (Zhao et al., 2024a), extracting context-aware guidelines (Fu et al., 2024), or checking belief states with fixed-question reflection (Kim et al., 2024).
@@
However, recent studies are depend on single-granularity memory derived from dynamic environmental interactions, whose effectiveness is fundamentally limited by the quality of the collected experiences. This constraint consequently reduces the variety of knowledge and limits the adaptability of planning. These limitations lead to explicit outcomes manifested in three aspects: (1) When confronted with challenging planning tasks, the initial exploration efficiency of LLMs is often suboptimal, thereby constraining the quality of memory. (2) Current approaches extract knowledge either at a holistic level or a scenario-specific level, both of which lead to homogenized granularity in the acquired knowledge. (3) Although directly using manually designed questions to reflect on trajectory information in short-term memory can yield decent results, it lacks generalizability across different environments.
@@
To bridge this gap, we introduce Coarse-to-Fine Grounded Memory (CFGM), a novel agent framework that leverages the internal knowledge of LLM to guide the memory collection and utilization, thereby enhancing their planning capabilities in complex environments. As illustrated in Figure 1, CFGM operates through three progressively memory grounding stages to ensure that the agent’s planning and actions remain continuously anchored in LLM’s prior understanding of memory. Guided by environmental information (initial memory), the LLM first identifies coarse-grained focus points for task planning. This preliminary understanding enables the collection of higher-quality experiences. For each one of the offline experiences (long-term memory), LLM grounds the knowledge it contains into high-level principles and detailed techniques, thereby constructing an experience-wise hybrid-grained tips dictionary. During online interactions, the most relevant successful experiences and tips are retrieved to enhance LLM planning as context. Upon facing environmental anomalies, the agent conducts a fine-grained key information analysis of the current state (short-term memory) , followed by a self-QA reflection grounded in both past successes and the present situation to effectively correct planning. By grounding explicit coarse-to-fine memories with LLM’s internal knowledge, CFGM enables the LLM to autonomously consider diverse information in complex planning environments and optimizes the collection and utilization of memory based on its reasoning outcomes. Because of this, CFGM is able to enhance the quality of collected memories, diversify experiential knowledge, and provide more flexible error-correction support during planning. Finally, this enables LLMs to tackle complex tasks more effectively, even preserving compatibility with powerful closed-source models. We evaluate CFGM on three diverse interactive planning benchmarks. Our results demonstrate that CFGM significantly outperforms existing memory-enhanced agent systems, showcasing improved performance and robustness.
@@
Our key contributions are as follows: (i) We propose Coarse-to-Fine Grounded Memory (CFGM), a novel framework that enhances LLM agents by systematically grounding memory with LLM’s internal knowledge during experiences collection, tips extraction and adaptive planning. (ii) We validate the effectiveness of CFGM across multiple challenging benchmarks, achieving state-of-the-art performance compared to strong memory-enhanced LLM agent system baselines. (iii) We perform comprehensive ablation studies and analyses on the framework’s core components, which show coarse-to-fine grounded memory enhances the quality of collected memory experiences, constructs hybrid-grained tips with more useful information, and enables flexible and effective reflection.
@@ 2.1 LLM-based Agents
As an emerging technology, LLM-based agents employ LLM which provides language understanding and generation capabilities as the core of planning (Yue, 2025). Compared to traditional AI agents, LLM-based agents demonstrate higher levels of autonomy and reasoning capabilities (Luo et al., 2025). LLM-based agents have been applied in various domains, including software engineering (Jin et al., 2024), question answering (Yue, 2025), scientific research (Xiao et al., 2024), web navigation (Deng et al., 2023), mathematical reasoning (Ahn et al., 2024), embodied intelligence (Zhao et al., 2024b), and search and recommendation systems (Zhang et al., 2025a). In this paper, we aim to construct high-performance agents targeting diverse planning task with coarse-to-fine grounded memory.
@@ 2.2 Experience-based Optimization for Agents
Experience-based optimization focuses on improving the performance of agents by leveraging historical data and accumulated knowledge (Li et al., 2024b, a; Zhao et al., 2024a; Chen et al., 2024; Fu et al., 2024; Qian et al., 2024). This optimization provides a mechanism for LLM agents to learn and adapt over time without adjusting the model’s parameters (Du et al., 2025; Zheng et al., 2025; Song et al., 2024). ExpeL (Zhao et al., 2024a) exemplifies this approach through prompt-based learning with experience recall and insight extraction. AutoGuide (Fu et al., 2024) extracts and applies context-aware guidelines from experiences. However, these methods often struggle with inefficient exploration in complex planning tasks due to limited prior task understanding, leading to suboptimal memory collection. Furthermore, when extracting knowledge from memories, they typically rely on single-grained information, restricting the agent’s holistic comprehension. To overcome these issues, CFGM leverages LLM-generated focus points to guide exploration toward critical dynamics, reducing redundancy. Additionally, it extracts hybrid-grained tips from experiences to enhance LLMs’ comprehensive memory understanding.
@@ 2.3 Feedback-based Optimization for Agents
Feedback-based optimization involves using various types of online feedback to guide the learning and improvement of agents. This feedback can come from external models (Bo et al., 2025; Huang et al., 2023; Yao et al., 2023b), or the agent itself (Huang et al., 2024; Wang et al., 2024b; Liang et al., 2024; Zhang et al., 2024; Zhou et al., 2024; Shinn et al., 2023b). Self-reflection and self-correction enable agents to iteratively improve their outputs by identifying and addressing errors (Luo et al., 2025). Reflexion (Shinn et al., 2023b) preforms trial-and-error attempts on online inference trajectories, reflects on the mistakes, and applies the lessons learned to subsequent attempts. QuBE  (Kim et al., 2024) conducts belief-state reflection with fixed-question templates, which addresses errors in tasks that allow only a single attempt. To enable more flexible short-term memory reflection in single-attempt planning compared to existing approaches, CFGM employs adaptive Self-QA reflection that dynamically leverage key information extracted from online trajectory information with LLM.
@@ 3 Method
To tackle the current challenges that memory-augmented agents encounter during planning, we introduce an LLM-based agent system characterized by a coarse-to-fine grounding of the model’s internal knowledge into explicit memories. This system leverages inherent LLM knowledge to enhance the quality of collected memories, diversify high-level experiential knowledge, and provide more flexible error-correction support during planning, ultimately enabling LLMs to handle complex decision-making tasks more effectively. The framework is depicted in Figure 2 and comprises three components: coarse-grained focus-driven experience collection, hybrid-grained experience-wise tips extraction and fine-grained key information adaptive planning. The prompt template related to memory grounding in the entire framework are presented in Appendix C.
Algorithm 1 Coarse-Grained Focus-Driven Experience Collection.
@@
4:  $t_{n}\leftarrow\mathcal{T_{\text{train}}}[n],\nu_{n,0}\leftarrow\text{\textquotedblleft\textquotedblright}$
@@
3.1 Coarse-Grained Focus-Driven Experience Collection
To collect usable and diverse experiential memories, we adopted a synergistic approach Reflexion Shinn et al. (2023a) for collaborative thinking and retry-reflection to collect diverse experiences. Additionally, we enabled the LLM to ground environmental foundational information to obtain focus points that guide LLM with their own preliminary understanding of the environment to collect higher-quality experience. The workflow of this component is illustrated in the pseudo-code of Algorithm˜1. Specifically, at the initiation of experience collection, the agent $\text{LLM}_{\text{Focus}}$ conducts a comprehensive analysis of the environmental description $Desc_{\text{env}}$ and manual trajectories few-shot $F_{\text{manual}}$ using its internal knowledge. Through coarse-grained examination of task-related foundational information, the agent can generate some preliminary thoughts for handling the environmental tasks. Guided by these focus points $FP$ on training task sets $\mathcal{T_{\text{train}}}$, the agent $\text{LLM}_{\text{ReAct}}$ executes up to $Z+1$ rounds of trial attempts involving think-act iterations. During this process, failures trigger introspective reflection $\text{LLM}_{\text{Reflect}}(\tau_{n,z})$, with updates incorporated into task-specific reflections $\nu_{n,z+1}$ to assist subsequent attempts until task $t_{n}$ completion. High-quality trajectory information $\tau_{n,z}$ collected under focus points guidance is stored in the experience pool $\mathcal{B}$ with the corresponding task $t_{n}$. Both successful and failed trajectories in success task will subsequently support tip generation and key information reflection phases.
Algorithm 2 Hybrid-Grained Experience-Wise Tips Extraction.
@@ 3.1 Coarse-Grained Focus-Driven Experience Collection
3.2 Hybrid-Grained Experience-Wise Tips Extraction
Because the experience collection stage follows a repeated trial-and-error paradigm, many tasks stored in the experience pool are accompanied by both successful and failed trajectories. These trajectories contain not only fine-grained scene-specific details but also more high-level coarse-grained insights that are applicable across similar tasks. We refer to such hybrid-grained knowledge as tips. To help the LLM make better use of this hybrid-grained information, the current module leverages the model’s internal knowledge offline to convert the multiple trajectories of a given task experience into a set of tips that blend hybrid-grained perspectives.
@@ 3.2 Hybrid-Grained Experience-Wise Tips Extraction
3.3 Fine-Grained Trajectory Information Adaptive planning
@@
The detailed workflow of the aforementioned evaluation reasoning is presented in Algorithm˜3. For task $t_{m}$ from evaluation set $\mathcal{T_{\text{evaluation}}}$, we conduct a prior analysis based on task description retrieval before inference. Using the Faiss library Johnson et al. (2019), we retrieve the top-k most similar experiences $E_{\text{sim}}$ from the Experience Pool $\mathcal{B}$ with textual embedder $\mathcal{E}$. These experiences’ successful trajectory contents $\tau_{h}^{\text{success}}$ and corresponding tips knowledge $TD[t_{h}]$ are integrated into Similar Trajectories content $ST$ and relevant Experience Tips $ET$ , which serve as contextual assistance for the decision-making of policy model $\text{LLM}_{\text{CFGM{}}}$ during inference. During the reasoning process for task $t_{m}$, we implement real-time monitoring of observed information $o_{i+1}$ and trigger alerts for specific abnormal observations. When triggered, we employ the Key Information Extraction model $\text{LLM}_{\text{KIE}}$ to perform a structured, fine-grained analysis of the current trajectory content $\tau_{m}$, thereby obtaining grounded current trajectory content as key information $KI$. Subsequently, by synthesizing $KI$, $\tau_{m}$, and $ST$ as contextual elements, we further instruct the Key Information Reflection model $\text{LLM}_{\text{KIR}}$ to conduct self-reflective question-answer $qa_{i}$, ultimately deriving a final corrective plan $ref_{i}$. The $ref_{i}$ is then integrated into the current trajectory content to guide $\text{LLM}_{\text{CFGM{}}}$ in implementing adaptive responses to exceptional situations.
Algorithm 3 Fine-Grained Trajectory Information Adaptive Planning.
@@
6:  $ET\leftarrow\text{\textquotedblleft\textquotedblright},ST\leftarrow\text{\textquotedblleft\textquotedblright}$
@@ 4.1 Experimental Settings
We evaluate CFGM on three challenging interactive environment to assess its effectiveness and capabilities, particularly how its coarse-to-fine grounding mechanism enhance LLM agent performance. We refer Appendix A for detailed information of main environments and WebArena (Zhou et al., 2023) used in ablation study.
Environments. The basic introduction of environment is as follows: (i) AlfWorld is an virtual household agent benchmark where the agent performs household tasks based on natural language instructions. It requires navigation and object manipulation (Shridhar et al., 2020). (ii) WebShop is an interactive environment simulating online shopping, where the agent needs to find and purchase a product matching specific criteria by interacting with a website (Yao et al., 2022). (iii) ScienceWorld is a text-based environment focusing on elementary science tasks, requiring complex reasoning and interaction with the simulated world (Wang et al., 2022).
@@
Baselines. We compare CFGM against some mainstream LLM agent methods, most of which are memory-augmented agents: (i) ReAct is a foundational method that interleaves reasoning steps with actions in the environment (Yao et al., 2023a). (ii) ExpeL learns from past experiences by extracting insights and retrieving relevant trajectories (Zhao et al., 2024a). (iii) AutoGuide generates context-aware guidelines from offline experiences to guide the agent at test time (Fu et al., 2024). (iiii) QuBE enhances agent reasoning by constructing belief state reflection with human-curated questions to alleviate reasoning derailment (Kim et al., 2024).

### v2

@@ 1 Introduction
Figure 1: Conceptual overview of CFGM. During the offline training, our method first extracts coarse-grained focus points to guide experience collection, then distills hybrid-grained tips from these experiences. At the online inference time, it retrieves relevant experiences and tips for planning. When encountering anomalies, the system identifies fine-grained key details for adaptive self-QA and plan adjustment.
@@
The remarkable progress of LLMs has ignited intense interest in their deployment as autonomous agents capable of tackling complex, long-horizon tasks within interactive environments (Wang2023; Xi2023; Gur2024; Wang2024). These LLM-based agents leverage extensive world knowledge (Zhao2023) and reasoning capabilities to formulate policies (yao2023react). However, significant challenges hinder their practical effectiveness. Agents often suffer from inefficient exploration, yielding sparse or irrelevant data (Arumugam2025; zhang2025beta; zhang2025edge). They also struggle with reasoning derailment in partially observable environments, where incomplete observations cause cascading errors (Kim et al., 2024). Existing methods attempt to mitigate these issues using memory mechanisms, such as storing offline experiences and insights (Zhao2024), extracting context-aware guidelines (Fu2024), or checking belief states with fixed-question reflection (Kim et al., 2024).
@@
However, recent studies depend on single-granularity memory derived from dynamic environmental interactions, whose effectiveness is limited by the quality of the collected experiences. This constraint consequently reduces the variety of knowledge and limits the adaptability of planning. These limitations lead to explicit outcomes manifested in three aspects: (1) When confronted with challenging planning tasks, the initial exploration efficiency of LLMs is often suboptimal, constraining the quality of memory. (2) Current approaches extract knowledge either at a holistic level or a scenario-specific level, both of which lead to homogenized granularity in the acquired knowledge. (3) Although directly using manually designed questions to reflect on trajectory information in short-term memory can yield decent results, it lacks generalizability across environments.
@@
To bridge this gap, we introduce Coarse-to-Fine Grounded Memory (CFGM), a novel agent framework that leverages the internal knowledge of LLM to guide the memory collection and utilization, thereby enhancing their planning capabilities in complex environments. As illustrated in Figure 1, CFGM operates through three progressively memory grounding stages to ensure that the agent’s planning and actions remain continuously anchored in LLM’s prior understanding of memory. Guided by environmental information (initial memory), the LLM first identifies coarse-grained focus points for task planning. This preliminary understanding enables the collection of higher-quality experiences. For each one of the offline experiences (long-term memory), LLM grounds the knowledge it contains into high-level principles and detailed techniques, thereby constructing an experience-wise hybrid-grained tips dictionary. During online interactions, the most relevant successful experiences and tips are retrieved to enhance LLM planning as context. Upon facing environmental anomalies, the agent conducts a fine‑grained key information analysis of the current state (short-term memory) , followed by a self-QA reflection grounded in both past successes and the present situation to effectively correct planning. By grounding explicit coarse-to-fine memories with LLM’s internal knowledge, CFGM enables the LLM to autonomously consider diverse information in complex planning environments and optimizes the collection and utilization of memory based on its reasoning outcomes. Because of this, CFGM is able to enhance the quality of collected memories, diversify experiential knowledge, and provide more flexible error‑correction support during planning. Finally, this enables LLMs to tackle complex tasks more effectively, even preserving compatibility with powerful closed‑source models. We evaluate CFGM on three diverse interactive planning benchmarks. Our results demonstrate that CFGM significantly outperforms existing memory-enhanced agent systems, showcasing improved performance and robustness.
@@
Our three key contributions are as follows: (i) We propose Coarse-to-Fine Grounded Memory (CFGM), a novel framework that enhances LLM agents by systematically grounding memory with LLM’s internal knowledge during experiences collection, tips extraction and adaptive planning. (ii) We validate the effectiveness of CFGM across multiple challenging benchmarks, achieving state-of-the-art performance compared to strong memory-enhanced LLM agent system baselines. (iii) We perform comprehensive ablation studies and analyses on the framework’s core components, which show coarse-to-fine grounded memory enhances the quality of collected memory experiences, constructs hybrid-grained tips with more useful information in offline phase, and enables flexible and effective reflection during online planning.
@@ 2.1 LLM-based Agents
As an emerging technology, LLM-based agents employ LLM which provides language understanding and generation capabilities as the core of planning (Yue2025). Compared to traditional AI agents, LLM-based agents demonstrate higher levels of autonomy and reasoning capabilities (Luo2025). LLM-based agents have been applied in various domains, including software engineering (Jin2024), question answering (Yue2025), scientific research (Xiao2024), web navigation (Deng2023), mathematical reasoning (Ahn et al., 2024), embodied intelligence (Zhao et al., 2024), and search and recommendation systems (zhang2025survey). In this paper, we aim to construct high-performance agents targeting diverse planning task with coarse-to-fine grounded memory.
@@ 2.2 Experience-based Optimization for Agents
Experience-based optimization focuses on improving the performance of agents by leveraging historical data and accumulated knowledge (Li2024g; Li2024c; Zhao2024; bai2024bayesian; Chen2024c; Fu2024; Qian et al., 2024; Fu et al., 2024; Zhang et al., ). This optimization provides a mechanism for LLM agents to learn and adapt over time without adjusting the model’s parameters (Du2025; Zheng2025; Song et al., 2024). ExpeL (Zhao2024) exemplifies this approach through prompt-based learning with experience recall and insight extraction. AutoGuide (Fu2024) extracts and applies context-aware guidelines from experiences. However, these methods often struggle with inefficient exploration in complex planning tasks due to limited prior task understanding, leading to suboptimal memory collection. Furthermore, when extracting knowledge from memories, they typically rely on single-grained information, restricting the agent’s holistic comprehension. To overcome these issues, CFGM leverages LLM-generated focus points to guide exploration toward critical dynamics, reducing redundancy. Additionally, it extracts hybrid-grained tips from experiences to enhance LLMs’ comprehensive memory understanding.
@@ 2.3 Feedback-based Optimization for Agents
Feedback-based optimization involves using various types of online feedback to guide the learning and improvement of agents. This feedback can come from external models (Bo2025; Huang2023; Yao2023; zhang2024exploiting; bai2024efficient; sun2025retrieval), or the agent itself (Huang et al., 2024; Wang et al., 2024; Liang2024c; Zhang et al., 2024; Zhou2024a; Shinn2023; cui2025self). Self-reflection and self-correction enable agents to iteratively improve their outputs by identifying and addressing errors (Luo2025). Reflexion (Shinn2023) preforms trial-and-error attempts on online inference trajectories, reflects on the mistakes, and applies the lessons learned to subsequent attempts. QuBE  (Kim et al., 2024) conducts belief-state reflection with fixed-question templates, which addresses errors in tasks that allow only a single attempt. To enable more flexible short-term memory reflection in single-attempt planning compared to existing approaches, CFGM employs adaptive Self-QA reflection that dynamically leverage key information extracted from online trajectory information with LLM.
@@ 3 Method
To tackle the current challenges that memory-augmented agents encounter during planning, we introduce an LLM‑based agent system characterized by a coarse‑to‑fine grounding of the model’s internal knowledge into explicit memories. This system leverages inherent LLM knowledge to enhance the quality of collected memories, diversify high‑level experiential knowledge, and provide more flexible error‑correction support during planning, ultimately enabling LLMs to handle complex decision‑making tasks more effectively. The framework is depicted in Figure 2 and comprises three components: coarse‑grained focus‑driven experience collection, hybrid‑grained experience‑wise tips extraction and fine‑grained key information adaptive planning. The prompt template related to memory grounding in the entire framework are presented in Appendix C.
Algorithm 1 Coarse‑Grained Focus‑Driven Experience Collection.
@@
4:  $t_{n}\leftarrow\mathcal{T_{\text{train}}}[n],\nu_{n,0}\leftarrow\text{“”}$
@@
3.1 Coarse‑Grained Focus‑Driven Experience Collection
To collect usable and diverse experiential memories, we adopted a synergistic approach Reflexion Shinn2023 for collaborative thinking and retry-reflection to collect diverse experiences. Additionally, we enabled the LLM to ground environmental foundational information to obtain focus points that guide LLM with their own preliminary understanding of the environment to collect higher-quality experience. The workflow of this component is illustrated in the pseudo-code of Algorithm˜1. Specifically, at the initiation of experience collection, the agent $\text{LLM}_{\text{Focus}}$ conducts a comprehensive analysis of the environmental description $Desc_{\text{env}}$ and manual trajectories few-shot $F_{\text{manual}}$ using its internal knowledge. Through coarse-grained examination of task-related foundational information, the agent can generate some preliminary thoughts for handling the environmental tasks. Guided by these focus points $FP$ on training task sets $\mathcal{T_{\text{train}}}$, the agent $\text{LLM}_{\text{ReAct}}$ executes up to $Z+1$ rounds of trial attempts involving think-act iterations. During this process, failures trigger introspective reflection $\text{LLM}_{\text{Reflect}}(\tau_{n,z})$, with updates incorporated into task-specific reflections $\nu_{n,z+1}$ to assist subsequent attempts until task $t_{n}$ completion. High-quality trajectory information $\tau_{n,z}$ collected under focus points guidance is stored in the experience pool $\mathcal{B}$ with the corresponding task $t_{n}$. Both successful and failed trajectories in success task will subsequently support tip generation and key information reflection phases.
Algorithm 2 Hybrid‑Grained Experience‑Wise Tips Extraction.
@@ 3.1 Coarse‑Grained Focus‑Driven Experience Collection
3.2 Hybrid‑Grained Experience‑Wise Tips Extraction
Because the experience collection stage follows a repeated trial-and-error paradigm (sutton2018reinforcement; dong2020deep), many tasks stored in the experience pool are accompanied by both successful and failed trajectories. These trajectories contain not only fine‑grained scene-specific details but also more high-level coarse‑grained insights that are applicable across similar tasks. We refer to such hybrid‑grained knowledge as tips. To help the LLM make better use of this hybrid-grained information, the current module leverages the model’s internal knowledge offline to convert the multiple trajectories of a given task experience into a set of tips that blend hybrid-grained perspectives.
@@ 3.2 Hybrid‑Grained Experience‑Wise Tips Extraction
3.3 Fine‑Grained Trajectory Information Adaptive planning
@@
The detailed workflow of the aforementioned evaluation reasoning is presented in Algorithm˜3. For task $t_{m}$ from evaluation set $\mathcal{T_{\text{evaluation}}}$, we conduct a prior analysis based on task description retrieval before inference. Using the Faiss library johnson2019billion, we retrieve the top-k most similar experiences $E_{\text{sim}}$ from the Experience Pool $\mathcal{B}$ with textual embedder $\mathcal{E}$. These experiences’ successful trajectory contents $\tau_{h}^{\text{success}}$ and corresponding tips knowledge $TD[t_{h}]$ are integrated into Similar Trajectories content $ST$ and relevant Experience Tips $ET$ , which serve as contextual assistance for the decision-making of policy model $\text{LLM}_{\text{CFGM{}}}$ during inference. During the reasoning process for task $t_{m}$, we implement real-time monitoring of observed information $o_{i+1}$ and trigger alerts for specific abnormal observations. When triggered, we employ the Key Information Extraction model $\text{LLM}_{\text{KIE}}$ to perform a structured, fine-grained analysis of the current trajectory content $\tau_{m}$, thereby obtaining grounded current trajectory content as key information $KI$. Subsequently, by synthesizing $KI$, $\tau_{m}$, and $ST$ as contextual elements, we further instruct the Key Information Reflection model $\text{LLM}_{\text{KIR}}$ to conduct self-reflective question-answer $qa_{i}$, ultimately deriving a final corrective plan $ref_{i}$. The $ref_{i}$ is then integrated into the current trajectory content to guide $\text{LLM}_{\text{CFGM{}}}$ in implementing adaptive responses to exceptional situations.
Algorithm 3 Fine‑Grained Trajectory Information Adaptive Planning.
@@
6:  $ET\leftarrow\text{“”},ST\leftarrow\text{“”}$
@@ 4.1 Experimental Settings
We evaluate CFGM on three challenging interactive environment to assess its effectiveness and capabilities, particularly how its coarse-to-fine grounding mechanism enhance LLM agent performance. We refer Appendix A for detailed information of main environments and WebArena (zhou2023webarena) used in ablation study.
Environments. The basic introduction of environment is as follows: (i) AlfWorld is an virtual household agent benchmark where the agent performs household tasks based on natural language instructions. It requires navigation and object manipulation (Shridhar2020). (ii) WebShop is an interactive environment simulating online shopping, where the agent needs to find and purchase a product matching specific criteria by interacting with a website (Yao2022). (iii) ScienceWorld is a text-based environment focusing on elementary science tasks, requiring complex reasoning and interaction with the simulated world (Wang et al., 2022).
@@
Baselines. We compare CFGM against some mainstream LLM agent methods, most of which are memory-augmented agents: (i) ReAct is a foundational method that interleaves reasoning steps with actions in the environment (yao2023react). (ii) ExpeL learns from past experiences by extracting insights and retrieving relevant trajectories (Zhao2024). (iii) AutoGuide generates context-aware guidelines from offline experiences to guide the agent at test time (Fu2024). (iiii) QuBE enhances agent reasoning by constructing belief state reflection with human-curated questions to alleviate reasoning derailment (Kim et al., 2024).